# The CLAS Collaboration

### v2

@@ Abstract
[…]
@@ I Introduction
Exclusive reactions of meson photo- and electroproduction off protons are intensively utilized in laboratories around the world as a very powerful tool to investigate nucleon structure and the principles of the strong interaction. These studies include the extraction of various observables from analyses of experimental data, as well as subsequent theoretical and phenomenological interpretations of the extracted observables Krusche and Schadmand (2003); Aznauryan and Burkert (2012); Skorodumina et al. (2015a); Ireland et al. (2020); Thiel et al. (2022).
Exclusive reactions off free protons have been studied in considerable detail, and a lot of information on differential cross sections and different single- and double-polarization asymmetries with almost complete coverage of the reaction phase space has become available. A large part of this information comes from the analysis of data collected in Hall B at Jefferson Lab (JLab) with the CLAS detector Mecking et al. (2003); CLAS Physics Database .
Meanwhile, reactions occurring in photon and electron scattering off nuclei have been less extensively investigated. Experimental information on these processes is sparse and mostly limited to inclusive measurements of total nuclear photoproduction cross sections Mokeev et al. (1995); Bianchi et al. (1994); Ahrens (1985) and the nucleon structure function $F_{\text{2}}$ Osipenko et al. (2005, 2006, 2010), while exclusive measurements off bound nucleons are lacking.
However, information on exclusive reactions off bound nucleons is crucially important to the investigation of nuclear structure and for a deeper understanding of the processes occurring in the nuclear medium because various exclusive channels will have different energy dependences and different sensitivities to the reaction mechanisms. This situation creates a strong demand for exclusive measurements off bound nucleons, and the deuteron, being the lightest and most weakly bound nucleus, is the best target for initiating these efforts.
@@
Exclusive reactions off bound protons manifest some specific features that are not present in reactions off free protons and originate from (a) Fermi motion of the initial proton and (b) final state interactions (FSIs) of the reaction final hadrons with the spectator neutron. In this paper special attention is paid to the detailed description of these issues.
[…]
@@
The details of the cross section extraction analysis are presented in Secs. III through VI, which encompass the selection of quasi-free events, the cross section calculation framework, the description of the corrections applied to the cross sections, and the study of the cross section uncertainties.
@@
Effects of the initial proton motion (also called Fermi motion) turned out to be tightly interwoven with many analysis aspects, and for this reason, their description is scattered throughout the paper. Meanwhile, FSI effects are addressed in a separate Section VII, which outlines specificities of FSIs in reactions off bound protons and their differences from FSIs in conventional free proton reactions. Some details on FSI effects in this particular analysis are also presented there.
@@
Section VIII presents the measured cross sections and their comparison with the cross section estimation obtained based on the JLab-MSU [^1] JLab - Moscow State University (Russia) model. model JM, which is a phenomenological reaction model for the process of double-pion production off free protons Mokeev et al. (2009, 2012, 2016).
@@
This study benefits from the fact that the free proton cross sections of the same exclusive reaction have been recently extracted from CLAS data Fedotov et al. (2017, 2018). These free proton measurements were performed under the same experimental conditions as in this study, including the beam energy value and the target setup. For this reason, the free proton study Fedotov et al. (2017, 2018), was naturally used as a reference point for many analysis components. This unique advantage allows one not only to verify the reliability of those analysis aspects that are similar for reactions off free and bound protons, but also to obtain a deeper understanding of those that differ. The latter include the effects of the initial proton motion and FSIs.
@@
Section IX introduces a comparison of the cross sections extracted in this study with their free proton counterparts from Refs. Fedotov et al. (2017, 2018), which allowed us to estimate the proportion of events in which FSIs between the final hadrons and the spectator neutron took place. Assuming the latter events to be the main cause of the difference between the cross section sets, their contribution to the total number of reaction events was found to vary in different regions of the reaction phase space, but in most parts of the phase space it maintains a level of $\sim$25%. Based on this comparison, different involved FSI mechanisms can further be investigated and other potential reasons that may contribute to the difference between the cross section sets can also be explored, which includes possible in-medium modifications of properties of nucleons and their excited states Mokeev et al. (1995); Bianchi et al. (1994); Ahrens (1985); Krusche (2005); Noble (1981).
@@ II Experiment
The “e1e” run period lasted from November 2002 until January 2003 and included several experiments with different beam energies (1 GeV and 2.039 GeV) and target cell contents (liquid hydrogen and liquid deuterium). This study is devoted to the experiment conducted with the 2-cm-long liquid-deuterium target and utilizing a 2.039-GeV electron beam.
@@ II.1 Detector setup
The magnet coils naturally separated the detector into six “sectors”, each functioning as an independent magnetic spectrometer. Each sector included four sub-detectors: drift chambers (DC), Čerenkov counters (CC), time-of-flight system (TOF), and electromagnetic calorimeter (EC) Mecking et al. (2003).
The azimuthal coverage for CLAS was limited only by the magnet coils and was approximately 90% at backward polar angles and 50% at forward angles Amarian et al. (2001). The polar angle coverage spanned from $8^{\circ}\mathrm{}$ to $45^{\circ}\mathrm{}$ for the Čerenkov counters and electromagnetic calorimeter and from $8^{\circ}\mathrm{}$ to $140^{\circ}\mathrm{}$ for the drift chambers and the time-of-flight system.
The drift chambers were located within the region of the magnetic field and performed charged particle tracking, allowing for the determination of the particle momentum from the curvature of their trajectories Mestayer et al. (2000). All other sub-detectors were located outside the magnetic field region, meaning that charged particles traveled through them along a straight line.
@@
The TOF scintillators were located radially outside the drift chambers and the Čerenkov counters but in front of the calorimeter. The time-of-flight system measured the time when a particle hit a TOF scintillator, thus allowing for the determination of the particle velocity. Then, using the particle momentum known from the DC, its mass can be determined, meaning that the particle can be identified Smith et al. (1999); Mutchler et al. (1998).
@@
The six CLAS sectors were equipped with a common data-acquisition (DAQ) system that collected the digitized data and stored the information for later off-line analysis.
@@ II.2 Target setup
The “e1e” target had a conical shape with the diameter varying from 0.35 to 0.6 cm, which served the purpose of effective extraction of gas bubbles formed in the liquid target content due to the heat that either originated from the beam and/or came from outside through the target walls. Due to the conical shape, the bubbles drained upwards and into a wider area of the target, thus clearing the beam interaction region and allowing the boiled deuterium to be effectively delivered back to the cooling system to be reliquified.
The target was located at --0.4 cm along the $z$ axis (near the center of CLAS) and its interaction region was 2-cm-long. The target cell had 15-$\mu$m-thick aluminum entrance and exit windows. In addition, an aluminum foil was located 2.0 cm downstream of the target. This foil was made exactly to the same specifications as the entry/exit windows of the target cell and served the purpose of providing reference event distributions for electron scattering off the target windows (see Sec. III.4.3).
More details on the “e1e” target assembly can be found in Ref. e1e target assembly .
@@ II.3 Data analysis strategy
Events corresponding to the investigated reaction $ep(n)\rightarrow e^{\prime}p^{\prime}(n^{\prime})\pi^{+}\pi^{-}$ were distinguished among all other detected events through the event selection procedure, described in detail in Sec. III. The selected exclusive events, however, represent only a part of the total number of events produced in the reaction, while the remainder were not registered due to (i) geometrical holes in the detector acceptance and (ii) less than 100% efficient particle detection within the detector acceptance. Therefore, to extract the reaction cross sections, the experimental event yield was adjusted for the geometric acceptance and detection efficiency, thereby accounting for the lost events.
In order to determine the overall detector efficiency, a Monte Carlo simulation was performed. In this analysis, double-pion events were generated with TWOPEG-D, which is an event generator for double-pion electroproduction off a moving proton Skorodumina et al. (2017a). These events are hereinafter called “generated” events.
@@
III Quasi-free event selection
For each event, the electron candidate was defined as the first-in-time particle that had signals in all four subcomponents of the CLAS detector (DC, CC, TOF, and EC). To select hadron candidates, signals only in two sub-detectors (DC and TOF) were required.
@@ III.1 Electron identification
To select good electrons among all electron candidates and to separate them from electronic noise, accidentals, and $\pi^{-}$ contamination, the electromagnetic calorimeter (EC) and Čerenkov counter (CC) responses were analyzed.
According to Ref. Egiyan et al. (1999), the overall EC resolution, as well as uncertainties from the EC output summing electronics, lead to fluctuations of the EC response near the hardware threshold. Therefore, to select only reliable EC signals, a minimal cut on the scattered electron momentum $P_{e^{\prime}}$ was applied in the software. The value of this cut was chosen to be 0.461 GeV, according to the relation suggested in Ref. Egiyan et al. (1999).
@@
In the next step, to eliminate part of the pion contamination, a sampling fraction cut was applied, which was based on different energy deposition patterns of electrons and pions in the EC. Specifically, when traveling through the EC, an electron produces an electromagnetic shower, where the deposited energy $E_{\text{tot}}$ is proportional to the electron momentum $P_{e^{\prime}}$, while a $\pi^{-}$ loses a constant amount of energy per scintillation layer independently of its momentum. Hence, for electrons the quantity $E_{\text{tot}}/P_{e^{\prime}}$ plotted as a function of $P_{e^{\prime}}$ is expected to follow a straight line parallel to the $x$-axis and located around the value 1/3 on the $y$-axis, since electrons lose about 2/3 of their energy in the EC lead sheets.
@@
Figure 2: Influence of different CC cuts on the photoelectron distributions for CLAS sector 3. Curves from top to bottom: black curve – only fiducial cut in the CC plane is applied, red curve – the $\varphi_{\text{cc}}$ matching cut is added, blue curve – the $\theta_{\text{cc}}$ matching cut is added, and green curve – the geometrical cut that removes inefficient CC zones is added. The peaking structure on the left is the contamination peak.
To further improve the quality of electron selection and $\pi^{-}/e^{-}$ separation, the Čerenkov counter response was analyzed Adams et al. (2001). Figure 2 illustrates photoelectron distributions measured in the CC for CLAS sector 3. As seen in Fig. 2, contamination is present in the measured CC spectra in the form of a peak located at values of a few photoelectrons. The contamination is thought to originate from accidental coincidences of photomultiplier tube (PMT) noise signals with measured pion tracks Osipenko et al. (2004). The goal of the event selection in the CC was to separate the spectrum of good electron candidates from the contamination peak, while minimizing the loss of good events. To achieve this goal, the following set of CC cuts was applied:
@@
The shape of the fiducial cut in the CC plane was taken from Ref. Khetarpal (2010). The $\varphi_{\text{cc}}$ and $\theta_{\text{cc}}$ matching procedures were based on the studies Osipenko et al. (2004) and Ungaro and Joo and relied on the anticipation that for real events, there must be a one-to-one correspondence between PMT signals in the CC and the angles in the CC plane (which are calculated from the DC information), while background noise and accidentals should not show such a correlation.
@@
The idea of the $\varphi_{\text{cc}}$ matching cut is that the particle track on the right side of the CC segment should match with the signal from the right-side PMT, and vice versa. Events that do not satisfy these conditions were removed. Events with signals from both PMTs were kept.
@@
Figure 3: $\theta_{\text{cc}}$ versus CC segment distributions for CLAS sector 2. Events between the horizontal black lines were treated as good electron candidates.
@@
To perform $\theta_{\text{cc}}$ matching, a $\theta_{\text{cc}}$ versus segment number cut was applied. Figure 3 shows the $\theta_{\text{cc}}$ versus segment distribution for CLAS sector 2. To develop the cut, event distributions in each segment were plotted as a function of $\theta_{\text{cc}}$ and fit with Gaussians. The horizontal black lines correspond to the positions of the fit maxima ±$\pm 4\sigma$. Events between these black lines were treated as good electron candidates Skorodumina et al. (2020); Skorodumina .
@@
Another important issue was that some specific geometrical zones in the CC showed low detection efficiency. When an electron hit such a zone, the number of detected photoelectrons was significantly less than expected. This leads to a systematic overpopulation of the low-lying part of the photoelectron spectrum and enhances the contamination peak. Since low efficiency zones were distributed inhomogeneously in the CC plane and the Monte Carlo simulation did not reproduce them properly, a geometrical cut that removes inefficient zones was applied. The details on this cut can be found in Refs. Skorodumina et al. (2020); Skorodumina ; Fedotov et al. (2017, 2018).
@@
The influence of the above cuts on the photoelectron distributions is demonstrated in Fig. 2, where the distribution before the matching cuts is plotted in black, after the $\varphi_{\text{cc}}$ matching – in red, and after the subsequent $\theta_{\text{cc}}$ matching cut – in blue. As seen in Fig. 2, the matching cuts reduce the contamination peak, but do not affect the main part of the photoelectron spectrum. Finally, the green distribution is plotted after adding the cut that removes inefficient CC zones. As expected, this cut leads to an event reduction in the low-lying part of the photoelectron spectrum, including the region of the contamination peak, while leaving the high-lying part of the spectrum essentially unchanged.
@@
The applied cuts result in a significant reduction of the contamination peak and its better separation from the main spectrum. As a final step, a cut on the number of photoelectrons was applied, which altogether eliminates the remains of the contamination peak. The cut position was individually optimized for each PMT for each CC segment for each CLAS sector. As the Monte Carlo did not reproduce photoelectron distributions well enough, the cut was performed only on the experimental data. To recover good electrons lost in this way, a standard procedure was applied, which was based on the fit of the photoelectron distributions by a modified Poisson function. More details on the procedure can be found in Refs. Skorodumina et al. (2020); Skorodumina ; Fedotov et al. (2017, 2018).

### v4

@@ Abstract
The single-differential and fully integrated cross sections for quasifree $\pi^{+}\pi^{-}$ electroproduction off protons bound in deuterium have been extracted for the first time. The experimental data were collected at Jefferson Laboratory with the CLAS detector. The measurements were performed in the kinematic region of the invariant mass $W$ from 1.3 to 1.825 GeV and the photon virtuality $Q^{2}$ from 0.4 to 1.0 GeV2. Sufficient experimental statistics allowed for narrow binning in all kinematic variables, while maintaining a small statistical uncertainty. The extracted cross sections were compared with the corresponding cross sections off free protons, which allowed us to obtain an estimate of the contribution from events in which interactions between the final-state hadrons and the spectator neutron took place.
@@ I Introduction
Exclusive reactions of meson photo- and electroproduction off protons are intensively studied in laboratories around the world as a very powerful tool to investigate nucleon structure and the principles of the strong interaction. These studies include extraction of various observables from analyses of experimental data, as well as subsequent theoretical and phenomenological interpretations of the extracted observables Krusche and Schadmand (2003); Aznauryan and Burkert (2012); Skorodumina et al. (2015a); Ireland et al. (2020); Thiel et al. (2022).
Exclusive reactions off free protons have already been studied in considerable detail, and a lot of information on differential cross sections and different single- and double-polarization asymmetries with almost complete coverage of the reaction phase space has become available. A large part of this information comes from analyses of data collected in Hall B at Jefferson Lab (JLab) with the CLAS detector Mecking et al. (2003); CLAS Physics Database .
Meanwhile, reactions occurring in photon and electron scattering off nuclei have been less extensively investigated. Experimental information on these processes is sparse and mostly limited to inclusive measurements of the total nuclear photoproduction cross section Mokeev et al. (1995); Bianchi et al. (1994); Ahrens (1985) and the nucleon structure function $F_{\text{2}}$ Osipenko et al. (2005, 2006, 2010), while exclusive measurements off bound nucleons are lacking.
However, information on exclusive reactions off bound nucleons is crucially important to the investigation of nuclear structure and for a deeper understanding of the processes occurring in the nuclear medium because various exclusive channels will have different energy dependences and different sensitivities to reaction mechanisms. This situation creates a strong demand for exclusive measurements off bound nucleons, and the deuteron, being the lightest and most weakly bound nucleus, is the best target with which to begin these efforts.
@@
Exclusive reactions off bound protons manifest some specific features that are not present in reactions off free protons and originate from (a) Fermi motion of the initial proton and (b) final-state interactions (FSIs) of the reaction final hadrons with the spectator neutron. In this paper, special attention is paid to the detailed description of these issues.
The paper introduces new information on the fully integrated and single-differential cross sections of the reaction $\gamma_{v}p(n)\rightarrow p^{\prime}(n^{\prime})\pi^{+}\pi^{-}$, where $\gamma_{v}$ denotes the virtual photon. The cross section measurements were performed in the kinematic region of the invariant mass $W$ from 1.3 to 1.825 GeV and the photon virtuality $Q^{2}$ from 0.4 to 1.0 GeV2. Sufficient experimental statistics allowed for narrow binning, i.e. 25 MeV in $W$ and 0.05 GeV2 in $Q^{2}$, while maintaining adequate statistical uncertainties. The extracted cross sections are quasifree, meaning that the admixture of events in which the final hadrons interacted with the spectator neutron is kinematically reduced to the achievable minimum.
@@
The details of the cross section extraction analysis are presented in Secs. III through VI, which encompass the selection of quasifree events, the cross section calculation framework, the description of the corrections applied to the cross sections, and the study of the cross section uncertainties, respectively.
@@
Effects of the initial proton motion (also called Fermi motion) turned out to be tightly interwoven with many analysis aspects, and for this reason their description is scattered throughout the paper. Meanwhile, FSI effects are addressed in separately in Sec. VII, which outlines specificities of FSIs in reactions off bound protons and their differences from FSIs in conventional free proton reactions. Some details on FSI effects in this particular analysis are also presented there.
@@
Section VIII presents the measured cross sections and their comparison with the cross section estimation obtained based on the JLab-MSU (JM) model [^1] JLab—Moscow State University (Russia) model., which is a phenomenological reaction model for the process of double-pion production off free protons Mokeev et al. (2009, 2012, 2016).
@@
This study benefits from the fact that the free proton cross sections of the same exclusive reaction have been recently extracted from CLAS data Fedotov et al. (2017, 2018). These free proton measurements were performed under the same experimental conditions as in this study, including the beam energy value and the target setup. For this reason, the free proton study Fedotov et al. (2017, 2018) was naturally used as a reference point for many analysis components. This unique advantage allowed us not only to verify the reliability of those analysis aspects that are similar for reactions off free and bound protons but also to obtain a deeper understanding of those that differ. The latter include the effects of the initial proton motion and FSIs.
@@
Section IX introduces a comparison of the cross sections extracted in this study with their free proton counterparts from Refs. Fedotov et al. (2017, 2018), which allowed us to estimate the proportion of events in which FSIs between the final hadrons and the spectator neutron took place. Assuming that events affected by FSIs with the spectator are the main cause of the difference between the cross section sets, their contribution to the total number of reaction events was found to be on a level of $\approx$25in the most part of the reaction phase space. With further theoretical interpretation of the performed comparison, various FSI mechanisms can be investigated and other potential reasons that may contribute to the difference between the cross section sets can be explored, which includes possible in-medium modifications of properties of nucleons and their excited states Mokeev et al. (1995); Bianchi et al. (1994); Ahrens (1985); Krusche (2005); Noble (1981).
@@ II Experiment
The “e1e” run period lasted from November 2002 until January 2003 and included several experiments with different beam energies (1 and 2.039 GeV) and target cell contents (liquid hydrogen and liquid deuterium). This study is devoted to the experiment conducted with the liquid-deuterium target and utilizing a 2.039-GeV electron beam.
@@ II.1 Detector setup
The magnet coils naturally separated the detector into six sectors, each functioning as an independent magnetic spectrometer. Each sector included four subsystems: drift chambers (DC), Čerenkov counters (CC), a time-of-flight (TOF) system, and an electromagnetic calorimeter (EC) Mecking et al. (2003).
The azimuthal coverage for CLAS was limited only by the magnet coils and was approximately 90% at backward polar angles and 50% at forward angles Amarian et al. (2001). The polar angle coverage spanned from $8^{\circ}\mathrm{}$ to $45^{\circ}\mathrm{}$ for the Čerenkov counters and the electromagnetic calorimeter and from $8^{\circ}\mathrm{}$ to $140^{\circ}\mathrm{}$ for the drift chambers and the time-of-flight system.
The drift chambers were located within the region of the magnetic field and performed charged particle tracking, allowing for determination of the particle momentum from the curvature of their trajectories Mestayer et al. (2000). All other subdetectors were located outside the magnetic field region, meaning that charged particles traveled through them along a straight line.
@@
The TOF scintillators were located radially outside the drift chambers and the Čerenkov counters but in front of the calorimeter. The time-of-flight system measured the time when a particle hit a TOF scintillator, thus allowing for determination of the particle velocity. Then, using the particle momentum known from the DC, its mass can be determined, meaning that the particle can be identified Smith et al. (1999); Mutchler et al. (1998).
@@
The six CLAS sectors were equipped with a common data-acquisition (DAQ) system that collected the digitized data and stored the information for later offline analysis.
@@ II.2 Target setup
The “e1e” target had a conical shape with the diameter varying from 0.35 to 0.6 cm, which served the purpose of effective extraction of gas bubbles formed in the liquid target contents due to the heat that either originated from the beam and/or came from outside through the target walls. Due to the conical shape, the bubbles drained upwards and into a wider area of the target, thus clearing the beam interaction region and allowing the boiled deuterium to be effectively delivered back to the cooling system to be reliquified.
The target was located at --0.4 cm along the $z$ axis (near the center of CLAS) and its interaction region was 2 cm long. The target cell had 15-$\mu$m-thick aluminum entrance and exit windows. In addition, an aluminum foil was located 2.0 cm downstream of the target. This foil was made exactly to the same specifications as the entry and exit windows of the target cell and served the purpose of providing reference event distributions for electron scattering off the target windows (see Sec. III.4.3).
In this experiment, the target setup was identical to that used for the free proton part of the “e1e” run period Fedotov et al. (2017, 2018). More details on the “e1e” target assembly can be found in Ref. e1e target assembly .
@@ II.3 Data analysis strategy
Events corresponding to the investigated reaction $ep(n)\rightarrow e^{\prime}p^{\prime}(n^{\prime})\pi^{+}\pi^{-}$ were distinguished among all other detected events through the event selection procedure described in detail in Sec. III. The selected exclusive events, however, represent only a part of the total number of events produced in the reaction, while the remainder were not registered due to (i) geometrical holes in the detector acceptance and (ii) less than 100% efficient particle detection within the detector acceptance. Therefore, to extract the reaction cross sections, the experimental event yield was adjusted for the geometric acceptance and detection efficiency, thereby accounting for the lost events.
To determine the overall detector efficiency, a Monte Carlo simulation was performed. In this analysis, double-pion events were generated with TWOPEG-D, which is an event generator for double-pion electroproduction off a moving proton Skorodumina et al. (2017a). These events are hereinafter called “generated” events.
@@
III Quasifree event selection
For each analyzed event, an electron candidate was defined as the first-in-time particle that had signals in all four subsystems of the CLAS detector (DC, CC, TOF, and EC). To select hadron candidates, signals only in two subdetectors (DC and TOF) were required.
@@ III.1 Electron identification
To select good electrons among all electron candidates and to separate them from electronic noise, accidentals, and $\pi^{-}$ contamination, various cuts in the electromagnetic calorimeter (EC) and the Čerenkov counters (CC) were applied.
According to Ref. Egiyan et al. (1999), the overall EC resolution, as well as uncertainties from the EC output summing electronics, led to fluctuations of the EC response near the hardware threshold. Therefore, to select only reliable EC signals, a minimal cut on the scattered electron momentum $P_{e^{\prime}}$ was applied in the software. The value of this cut was chosen to be 0.461 GeV, according to the relation suggested in Ref. Egiyan et al. (1999).
@@
In the next step, to eliminate part of the pion contamination, a sampling fraction cut was applied, which was based on different energy deposition patterns of electrons and pions in the EC. Specifically, when traveling through the EC, an electron produces an electromagnetic shower, where the deposited energy $E_{\text{tot}}$ is proportional to the electron momentum $P_{e^{\prime}}$, while a $\pi^{-}$ loses a constant amount of energy per scintillation layer independently of its momentum. Hence, for electrons, the quantity $E_{\text{tot}}/P_{e^{\prime}}$ plotted as a function of $P_{e^{\prime}}$ is expected to follow a straight line parallel to the $x$ axis and located around the value 1/3 on the $y$ axis (since electrons lose about 2/3 of their energy in the EC lead sheets).
@@
Figure 2: Influence of different CC cuts on the photoelectron spectrum for CLAS sector 3. Curves from top to bottom: the black curve – only the fiducial cut in the CC plane is applied, the red curve – the $\varphi_{\text{cc}}$ matching cut is added, the blue curve – the $\theta_{\text{cc}}$ matching cut is added, and the green curve – the geometrical cut that removes inefficient CC zones is added. The peaking structure on the left is the contamination peak.
To further improve the quality of electron identification and $\pi^{-}/e^{-}$ separation, event selection in the CC was performed Adams et al. (2001). Figure 2 illustrates photoelectron distributions measured in the CC for CLAS sector 3. As seen in Fig. 2, contamination is present in the CC spectrum in the form of a peak located at values of a few photoelectrons. The contamination is thought to originate from accidental coincidences of measured pion tracks with noise signals from photomultiplier tubes (PMTs) Osipenko et al. (2004). The goal of event selection in the CC was to separate the spectrum of good electron candidates from the contamination peak, while minimizing the loss of good events. To achieve this goal, the following set of CC cuts was applied:
@@
The analytical shape of the fiducial cut in the CC plane was taken from Ref. Khetarpal . The $\varphi_{\text{cc}}$ and $\theta_{\text{cc}}$ matching procedures were based on the studies Osipenko et al. (2004) and Ungaro and Joo and relied on the anticipation that for real events, there must be a one-to-one correspondence between PMT signals in the CC and the angles in the CC plane (which are calculated from the DC information), while background noise and accidentals should not show such a correlation.
@@
The idea of the $\varphi_{\text{cc}}$ matching cut is that the particle track on the right side of the CC segment should match with the signal from the right-side PMT, and vice versa. Events that did not satisfy these conditions were removed. Events with signals from both PMTs were kept.
@@
Figure 3: $\theta_{\text{cc}}$ versus CC segment number distribution for CLAS sector 2. Events between the horizontal black lines were treated as good electron candidates.
@@
To perform $\theta_{\text{cc}}$ matching, a $\theta_{\text{cc}}$ versus segment number cut was applied. Figure 3 shows the $\theta_{\text{cc}}$ versus segment number distribution for CLAS sector 2. To develop the cut, event distributions in each segment were plotted as a function of $\theta_{\text{cc}}$ and fit with Gaussians. The horizontal black lines correspond to the positions of the fit maxima ±$\pm 4\sigma$. Events between these black lines were treated as good electron candidates Skorodumina et al. (2020); Skorodumina .
@@
Another important issue was that some specific geometrical zones in the CC showed low detection efficiency. When an electron hit such a zone, the number of detected photoelectrons was significantly less than expected. This led to a systematic overpopulation of the low-lying part of the photoelectron spectrum and an enhancement of the contamination peak. Low-efficiency zones were distributed inhomogeneously in the CC plane and (being too dependent on specific features of the CC design) could not be simulated by the Monte Carlo technique. For this reason, a geometrical cut that removes inefficient zones was applied to both experimental events and reconstructed Monte Carlo events. The details on this cut can be found in Refs. Skorodumina et al. (2020); Skorodumina ; Fedotov et al. (2017, 2018).
@@
The influence of the above cuts on the photoelectron spectrum is demonstrated in Fig. 2, where the distribution before the matching cuts is plotted in black, after the $\varphi_{\text{cc}}$ matching – in red, and after the subsequent $\theta_{\text{cc}}$ matching cut – in blue. As seen in Fig. 2, the matching cuts reduce the contamination peak but do not affect the main part of the photoelectron spectrum. Finally, the green distribution is plotted after adding the cut that removes inefficient CC zones. As expected, this cut leads to an event reduction in the low-lying part of the photoelectron spectrum (including the region of the contamination peak), while leaving the high-lying part of the distribution essentially unchanged.
@@
The applied cuts result in a significant reduction of the contamination peak and its better separation from the main part of the photoelectron spectrum. As a final step, an explicit cut on the number of photoelectrons was applied, which altogether eliminated the remains of the contamination peak. The cut position was individually optimized for each PMT for each CC segment for each CLAS sector. As the Monte Carlo simulation did not reproduce photoelectron distributions well enough, the cut was performed only on the experimental data. To recover good electrons lost in this way, a standard correction procedure was applied, which was based on the fit of the photoelectron distributions by a modified Poisson function. More details on the procedure can be found in Refs. Skorodumina et al. (2020); Skorodumina ; Fedotov et al. (2017, 2018).